# AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library

### v2

@@ 3.1.1 Library Learning
The objective of this stage is to generate reusable insights as structured 4-tuples (Taxonomy, Condition, Explanation, Example) and organize them in a hierarchical taxonomy for efficient retrieval, while minimizing redundancy in the library. The left panel of Figure 3 (see the Library learning flow label) illustrates the workflow for this stage.
Insight Extraction, Representation, and Supervision Mode.
@@
Insights can be learned from either problems with a gold-standard program or with the answer alone. For each task, the system first constructs a mathematical formulation, then generates an executable solver program, and invokes the solver. When the library is non-empty, both steps are guided by retrieved insights. If the generated program does not achieve the correct optimal value, two supervision modes are used to guide the generation of new insights. When a gold program is available, the system compares the candidate formulation and program against the reference, diagnosing discrepancies (e.g., missing variables, misformulated constraints, incorrect objective terms) and distilling them into insights. When only the answer (i.e., final optimal objective) is provided, the system performs solver-guided self-exploration: it iteratively proposes executable programs, reuses prior failures as context, and receives verification from the solver. Once a program achieves its correct objective, it is treated as a proxy for the gold standard in anchor insight extraction. Before being stored in the library, each insight is locally verified by reapplying it to its source task to ensure that it resolves the original failure. In addition, to mitigate stochastic successes that could obscure useful lessons, we conduct three independent trials per task, allowing errors from probabilistic generation to serve as signals for learning.
@@
Each insight is represented as a structured 4-tuple: Taxonomy, hierarchical labels for indexing and retrieval; Condition, an explicit description of the applicability signals in the problem; Explanation, the underlying principle of applying this insight; and Example, a concrete demonstration such as a mathematical constraint or code snippet.
Library Storage and Retrieval.
@@
Insights are stored in a dynamically updated hierarchical taxonomy organized into three main tracks: Domain Modeling (problem-specific structures and assumptions), General Formulation (reusable mathematical patterns), and Code Implementation (solver-specific coding practices). Under each track, insights are further classified with two-level labels, where Level-1 captures a broad category and Level-2 refines it into a more specific subcategory. The taxonomy is initialized with few-shot labels and expands online: each new insight is either mapped to an existing category or, if no suitable label exists, prompts the LLM to propose new Level-1 or Level-2 labels. Each label is also assigned a condition, written by the LLM, that specifies when the category should be retrieved. When storing insights, to reduce redundancy, the LLM also checks whether a similar insight already exists and performs merging when appropriate. To align a target task with relevant insights, we employ a two-step LLM-driven retrieval procedure: Quick label matching, then full applicability check. The system first scans the taxonomy dictionary to identify labels that are potentially relevant to the context of the tasks. For example, Level-2 label such as Fixed Charge (Big-M Linking) will be probably detected when the problem description specifies that service or flow is allowed only if a facility is opened. After candidate labels are identified, the system rigorously evaluates each associated insight by examining its condition, and only the most applicable insights are retained.
During solution generation, retrieved insights from the Domain Modeling and General Formulation tracks guide the construction of the mathematical model, while insights from the Code Implementation track guide solver-code generation. This two-step procedure ensures that insights are applied appropriately, while the hierarchical taxonomy provides an extensible structure for organizing and retrieving insights as the library grows.
Operational Flow.
Training proceeds in an online regime over minibatches of data, starting from an empty library. For each batch, the system retrieves candidate insights, generates and executes programs, and upon failures extracts insights and immediately commits those that pass local self-verification, allowing newly added insights to benefit subsequent tasks and preventing the generation of repetitive insights. To reduce redundancy, tasks are clustered and reordered by problem type and semantic similarity, and overlapping insights within a batch are merged prior to integration. The process iterates until overall accuracy plateaus, at which point the library is archived and used for evaluation.
In implementation, for the sake of efficiency, training follows two coordinated data flows. The first processes minibatches of tasks in parallel for insight extraction. The second maintains a centralized queue of all generated insights, storing them into the library sequentially. This queue does not allow asynchronous updates, as concurrent modifications could lead to conflicts if two insights attempt to update the library simultaneously. This design balances parallelism in problem-solving with strict serialization in library updates, ensuring both efficiency and consistency.
3.1.2 Library Evolution
Figure 4: A locally verified initial condition (light-grey dashed circle) is refined into a broader applicability boundary (solid black) through evolutionary prompt optimization guided by the aggregate metric.
@@ 3.1.2 Library Evolution
While Library Learning expands the repository of insights, Library Evolution aims to transform task-specific lessons into broadly applicable knowledge. Since each insight’s applicability is defined by a condition induced from a specific task, early conditions are often too narrow (failing to trigger on relevant tasks) or too broad (causing misretrieval). Left unchecked, these misalignments lead to missed opportunities or systematic failures. Library Evolution counters this with a diagnostic–refinement cycle: it detects misaligned insights, aggregates evidence across tasks, and refines conditions at the end of each iteration. The refinement is guided by an aggregate metric rather than ad-hox fixes. As illustrated in Figure 4, library refinement can be understood as adjusting each insight’s condition toward the correct retrieval boundary in the problem space.
Diagnostic: Library Diagnosis.
@@
After each training round, we trace failed tasks and analyze their interaction with the library. The diagnostic agent partitions the relationship between each insight $i$ and its associated tasks into three disjoint categories: $\Pi(i)=\{\text{Positive}:S_{i}^{+},\text{Negative}:S_{i}^{-},\text{Unretrieved}:S_{i}^{u}\}$ where $S_{i}^{+}$ contains tasks where the insight was applicable and contributed to the correct formulation, $S_{i}^{-}$ contains tasks where it was misleading and degraded performance, and $S_{i}^{u}$ contains tasks where it was not retrieved but would have been beneficial. By maintaining these partitions across iterations, the system continuously builds a performance profile for each insight. If a failed task is subsequently solved after removing a misleading (negative) insight or by injecting a previously unretrieved one, the system attributes the failure to condition misalignment rather than lack of knowledge, thus avoiding redundant insight generation. Unretrieved tasks are identified by first comparing the model’s generated program with the ground-truth (or a reference program from self-exploration) to locate discrepancies. These discrepancies guide the search for candidate insights, which are then verified for their ability to resolve the errors. Verified insights are flagged as relevant but unretrieved, allowing the system to diagnose retrieval gaps without resorting to intractable combinatorial search. The left panel of Figure 3 (see the Library diagnosis flow label) illustrates the workflow for this stage.
Evolver: Library Refinement.
@@
Building on the diagnosis, the Evolver agent refines insights in two steps: condition refinement and refinement verification. First, for each diagnosed insight, the agent strengthens or prunes its applicability condition. Negative tasks contribute explicit inapplicability clauses (e.g., constraints or contexts that block use), while unretrieved tasks highlight missing applicability signals. The Evolver then proposes multiple refinement strategies (e.g., adding preconditions, introducing keyword anchors, merging overlapping triggers) and produces candidate conditions with the goal of preserving correct cases, eliminating mismatches, and recovering previously missed tasks. Then, each candidate condition replaces the original and is tested over the union $R_{i}=S_{i}^{+}\cup S_{i}^{-}\cup S_{i}^{u}$. A performance score
@@
quantifies improvement. Here, “kept positives” are tasks that still remain correctly retrieved after refinement; “corrected negatives” are tasks that were misled by the insight before and no longer retrieved; and “recovered unretrieved” are tasks that become correctly retrieved after refinement. We accept refinements that increase $p_{i}$ and keep the one with the highest $p_{i}$. The workflow for this stage is illustrated in the right panel of Figure 3.
3.2 Optimization Perspective
@@
$\max_{\ell\in\mathcal{L}}\;\;\mathbb{E}_{t\sim\mathcal{T}}\!\left[\mathrm{Success}(t\mid\ell)\right]\;-\;\lambda\,\Omega(\ell).$
@@
where $\mathrm{Success}(t\mid\ell)$ indicates whether $\ell$ enables the system to produce a program that achieves the correct optimal objective for task $t$, and $\Omega(\ell)$ quantifies library complexity (e.g., number of insights or redundancy-adjusted size). Under our problem design—bounded and continuous property of $\text{Success}(\cdot)$ and $\Omega(\cdot)$, sufficient exploration under solver verification, and bounded merging—the refinement dynamics converge to a locally optimal library. In Appendix C, we provide a conceptual sketch showing that convergence holds: As refinement in the second phase strictly improves the first term, while verified merging in the first phase reduces the second term without diminishing the first, sufficient exploration combined with iterative cycles of library learning and evolution ensures convergence to a local optimum. Given the inherent ambiguity of natural language and stochasticity in LLM outputs, we present this perspective not as a strict theorem but as a principled justification for the acquisition–refinement design and the redundancy-reduction operations.
4 EXPERIMENTS
@@
Our experiments are designed to reflect the requirements that arise in real-world optimization and operations research (OR) applications. In these settings, methods are expected not only to perform well on standard benchmarks, but also to transfer across domains, to remain effective when limited supervision is available, to improve steadily as more data becomes available, and to offer results that can be inspected and audited. We therefore organize our evaluation around four questions: (1) How well does the method generalize across domains? (2) Can it learn effectively with limited supervision? (3) Does performance improve consistently with more training data? (4) How does it compare overall with strong baselines? Finally, we examine the interpretability of the insight library to assess whether the outputs are understandable and actionable to practitioners.
4.1 Experimental Setup
@@
Our experiments are conducted on a dataset of 454 problem instances, aggregated from four real-world optimization and operation task datasets, namely the NLP4LP (AhmadiTeshnizi et al., 2024), NL4OPT (Ramamonjison et al., 2022), IndustryOR (Huang et al., 2025), MAMO (ComplexLP) (Huang et al., 2024), with any invalid entries discarded. These collections span various formulation types and originate from diverse sources, including academic papers, textbooks, and real‐world industry scenarios. Detailed descriptions of these datasets are provided in Appendix B.
@@
We perform stratified sampling within each dataset, randomly partitioning $70\%$ for training and $30\%$ for testing. We maintain a strict separation between training and test data. The experience library is constructed only from training tasks, ensuring that training-derived insight examples do not leak into the test set. To assess out-of-distribution generalization, we additionally evaluate on LogiOR (Yang et al., 2025a) and Optibench (Yang et al., 2024).
@@
Unless otherwise specified, GPT-4o (OpenAI 2024) with temperature 0 is used as the backbone. We use the success rate as the primary evaluation metric, following the evaluation protocol of Yang et al. (2025a), where a task is considered successful if the LLM-generated optimal value closely aligns with the provided ground-truth solution.
Baselines. We evaluate against two families of baselines. (i) Prompt-based: a vanilla baseline that directly generates the mathematical model from a simple prompt, as well as Reflexion (Shinn et al., 2023), OptiMUS (AhmadiTeshnizi, Gao, and Udell, 2024), and ORThought (Yang et al., 2025a). (ii) Learning-based: ORLM (Huang et al., 2025), built on LLaMa3-8B, and LLMOPT (Jiang et al., 2025), built on Qwen2.5-14B (the latest open-source version available after their paper).
Figure 5: Performance on out-of-distribution datasets. Numbers report test accuracy on LogiOR, OptiBench, and OptMath.
4.2 Out-of-Distribution Generalization
We evaluate how well different methods generalize beyond their training distribution. For this purpose, we use two benchmarks that were not included during training: LogiOR (Yang et al., 2025a) and OptiBench (Yang et al., 2024). Details are provided in Appendix B.
These datasets were either released after the baseline model (ORLM) or explicitly designed in baseline model’s experiment setting to avoid overlap with their training set (LLMOPT).
Figure 5 summarizes the results. Fine-tuned models such as ORLM and LLMOPT show strong in-distribution performance but exhibit a noticeable drop on unseen datasets. For example, ORLM falls to $19.6\%$ on LogiOR and $13.3\%$ on OptMath, while LLMOPT performs better but still degrades compared to its in-distribution performance. By contrast, AlphaOPT maintains higher accuracy across all three benchmarks, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench. These results support our hypothesis: fine-tuned models tend to learn the syntax of solutions and may perform well when problems are very similar, but they struggle to capture the underlying principles needed for broader problem solving. In contrast, the learned experience library equips AlphaOPT with stronger out-of-distribution generalization capability.
4.3 Learning with Limited Supervision
In practical OR applications, gold-standard programs are rarely available. We therefore test whether AlphaOPT can learn solely from answers. Since two datasets in our training set contain gold-standard programs, we remove them in this experiment and let AlphaOPT learn exclusively from answer labels through self-exploration, as introduced in Section 3.1.1. As shown in the last two rows of Table 2, remarkably, when trained with answer-only supervision, AlphaOPT achieves accuracy comparable to when it is trained with gold-standard programs. AlphaOPT (self-exploration) outperforms all prompt-based methods on test splits of the training data and even achieves the best performance on the OOD OptiBench dataset (92.1% accuracy). This demonstrates another advantage of AlphaOPT over fine-tuning–based methods, which require detailed annotations of mathematical formulations and code in order to achieve strong performance.
Table 1: AlphaOPT steadily improves in both Micro and Macro averages with increasing training size, while maintaining a compact library.
Training Size MicroAvg MacroAvg Library Size
100 83.24% 65.80% 38
200 85.09% 69.22% 103
300 85.21% 72.12% 110
4.4 Continual Growth with Data
We test whether AlphaOPT can improve its performance as more data becomes available. We incrementally sample sets of 100, 200, and 300 data items from our training set and train AlphaOPT on each subset. As shown in Table 1, when evaluated on out-of-distribution datasets (LogiOR, OptiBench), we observe that AlphaOPT steadily improves its performance with increasing data size, without requiring updates to its model parameters.
4.5 Overall Performance

### v3

@@ 3.1. Library Learning
This stage is designed to generate reusable insights as structured 4-tuples (Taxonomy, Condition, Explanation, Example) and organize them in a hierarchical taxonomy dictionary for efficient retrieval while minimizing redundancy in the library. Figure 2 (Library Learning Flow) clearly illustrates the overall workflow.
Insight Extraction and Representation
@@
For each arriving task, the system first constructs a mathematical formulation, then generates an executable solver program and invokes the solver. When the library is non-empty, both steps are guided by retrieved insights. If the generated program does not achieve the correct optimal value, the system performs solver-guided self-exploration: It iteratively proposes executable programs, reuses prior failures as context, and receives verification from the solver. Once a program achieves its correct objective, it is treated as a proxy for the gold standard in anchor insight extraction.
@@
Each insight is represented as a structured 4-tuple: Taxonomy, hierarchical labels for indexing and retrieval; Condition, an explicit description of the applicable prerequisites and triggering signals in the problem; Explanation, the underlying principle of applying this insight; and Example, a concrete demonstration such as corresponding mathematical expressions or code snippet.
Insight Verification.
To ensure the retrievability and correctness of newly generated natural-language insights and to prevent the uncontrolled accumulation of idle or invalid entries in the library, each insight undergoes two stages of strict verification before being stored: Retrieval verification: temporarily incorporate the insight into a copy of the library, and check whether the source task can match it via LLM-driven retrieval; Execution verification: reapply the new insight together with other retrieved insights to its source task to verify that the new insight can resolve the original failure and lead to optimality.
At each stage, a limited number of rewrite attempts is allowed until the insight passes verification, during which the LLM performs targeted revisions of the taxonomy, condition or the whole insight to improve its efficacy.
Library Storage.
@@
Experience library is maintained as a dynamically updating taxonomy dictionary organized into three main tracks: Domain Modeling (problem-specific structures and assumptions), General Formulation (reusable mathematical patterns), and Code Implementation (solver-specific coding practices). Within each track, the library organizes insights using a two-level labeling scheme, where Level-1 captures a broad category and Level-2 specializes it into a more specific subcategory. The taxonomy dictionary is initialized with few-shot labels and expands online: Each new insight is either mapped to an existing category or, if no suitable label exists, prompts the LLM to propose new Level-1 or Level-2 labels. Each label is also assigned a condition, written by the LLM, that specifies when the category should be retrieved.
To prevent uncontrolled growth of the experience library, we incorporate an online merging mechanism throughout training to prevent degradation in retrieval and execution efficiency. For each newly generated insight, the system performs redundancy checking with existing insights, merging when appropriate, followed by correctness validation on all associated tasks.
Library Retrieval.
During solution generation, retrieved insights from the Domain Modeling and General Formulation tracks guide the construction of the mathematical model, while insights from the Code Implementation track guide solver-code generation. To align a target task with relevant insights, we employ a two-step LLM-driven retrieval procedure: Quick label matching, then full applicability check. The system first scans the taxonomy dictionary to identify labels that are potentially relevant to the context of the tasks. For example, a Level-2 label such as Fixed Charge (Big-M Linking) will probably be detected when the problem description specifies that service or flow is allowed only if a facility is opened. After candidate labels are identified, the system rigorously evaluates each associated insight by examining its condition, and only the most applicable insights are retained. By employing a progressive disclosure strategy that reveals only task-relevant information when it becomes applicable, this procedure mitigates context overflow while maintaining the effectiveness of retrieval.
3.2. Library Evolution
@@
While Library Learning expands the repository of insights, Library Evolution aims to transform task-specific lessons into broadly applicable knowledge. Since each insight’s applicability is defined by a condition induced by a specific task, early conditions are often too narrow (failing to trigger on relevant tasks) or too broad (causing misretrieval). Left unchecked, these misalignments lead to widespread misguidance or missed opportunities. Library Evolution counters this with a diagnosis–refinement cycle: It detects misaligned insights, aggregates evidence across tasks, and refines conditions at the end of each iteration. The refinement is guided by an aggregate metric rather than ad hoc fixes. As illustrated in Figure 3, library refinement can be understood as adjusting each insight’s condition toward the correct retrieval boundary in the problem space.
Library Diagnosis.
@@
After an initial pass over the training tasks, the system continuously diagnoses task failures and traces their interactions with library insights through iterative cycles of insight retrieval and solution generation across tasks. By comparing the formulation and program generated under the guidance of retrieved insights with the ground truth to locate discrepancies, the diagnostic LLM agent partitions the relationship between each insight $i$ and its associated tasks into three disjoint categories: $\Pi(i)=\{\text{Positive}:S_{i}^{+},\text{Negative}:S_{i}^{-},\text{Unretrieved}:S_{i}^{u}\}$ where $S_{i}^{+}$ contains tasks for which the insight was applicable and contributed to the correct formulation and program, $S_{i}^{-}$ contains tasks for which it was misleading and degraded performance, and $S_{i}^{u}$ contains tasks for which it was not retrieved but would have been beneficial. By maintaining these partitions across iterations, the system continuously builds a performance profile for each insight. If a failed task is subsequently solved after removing a misleading (negative) insight or by injecting a previously unretrieved one, the system attributes the failure to condition misalignment rather than lack of knowledge and thus avoids redundant insight generation. Figure 2 (Library Diagnosis Flow) illustrates the workflow for this stage.
Figure 3. A locally verified initial condition is refined into a broader applicability boundary through LLM-driven refinement.
The figure provides an abstract illustration of how an insight’s applicability condition is refined over time. Each point represents a task outcome, where green points indicate tasks for which the insight is helpful, red points indicate tasks for which it is harmful, and blue points indicate tasks where the insight was not retrieved. The dashed curve shows the initial applicability condition, which is locally verified but narrow. Through library refinement, the condition boundary is adjusted, shown as the solid curve, to better separate positive cases from negative ones and to cover relevant unretrieved cases. The arrow indicates the refinement process, resulting in a broader and more accurate applicability boundary for the insight.
Library Refinement.
@@
Building on the diagnosed interaction of each insight, the LLM agent first strengthens or prunes its applicability condition. Negative tasks contribute explicit inapplicability clauses (e.g., constraints or contexts that block use), while unretrieved tasks highlight missing applicability signals. Multiple refinement strategies are asked to be proposed (e.g., adding preconditions, introducing keyword anchors, merging overlapping triggers) and produce candidate conditions with the goal of preserving correct cases, eliminating mismatches, and recovering previously missed tasks. Then, each candidate condition replaces the original and is tested over the union $R_{i}=S_{i}^{+}\cup S_{i}^{-}\cup S_{i}^{u}$. A performance score
@@
quantifies improvement. Here, kept positives are tasks that remain correctly retrieved after refinement; corrected negatives are tasks that were misled by the insight before and are no longer retrieved; and recovered unretrieved are tasks that become correctly retrieved after refinement. We accept refinements that increase $p_{i}$ and keep the one with the highest $p_{i}$. The workflow for this stage is illustrated in the right panel of Figure 2.
Table 1. Accuracy (%) on Test Split and Out-of-Distribution datasets, along with micro- and macro-averaged accuracy across datasets. Best per column in bold and second best underlined (ablation variants excluded from highlighting).
Test Split Out-of-Distribution
Method NLP4LP (73) NL4OPT (64) IndustryOR (25) MAMO (ComplexLP) (34) Micro Macro LogiOR (92) OptiBench (403) Micro Macro
Prompt-based Standard 68.5 54.7 52.0 44.1 57.7 54.8 46.7 72.7 67.9 59.7
OptiMus 71.2 73.4 36.0 29.4 60.2 52.5 17.4 74.7 64.1 46.0
ORThought 69.9 75.0 60.0 41.2 65.3 61.5 44.6 84.4 77.0 64.5
Fine-tuning ORLM 86.3 87.5 36.0 55.9 75.0 66.4 19.6 78.2 67.3 48.9
LLMOPT 97.3 86.5 44.0 85.8 85.0 78.4 40.2 66.4 61.5 53.3
Experience-learning Reflexion 76.7 64.1 56.0 47.1 64.8 61.0 43.5 76.9 70.7 60.2
Expel 79.5 67.2 48.0 70.4 69.9 66.3 47.8 80.4 74.3 64.1
[…]
Ablation (AlphaOPT) w/o Refinement 82.2 ($\downarrow 4.1$) 80.7 ($\downarrow 0.5$) 44.0 ($\downarrow 16.0$) 61.8 ($\downarrow 14.7$) 73.3 ($\downarrow 6.8$) 67.2 ($\downarrow 9.2$) 48.9 ($\downarrow 7.1$) 92.1 ($\downarrow 1.4$) 84.1 ($\downarrow 2.4$) 70.5 ($\downarrow 4.3$)
w/o Insight Example 86.3 ($\downarrow 1.4$) 78.1 ($\downarrow 3.1$) 56.0 ($\downarrow 4.0$) 58.8 ($\downarrow 17.7$) 76.0 ($\downarrow 4.1$) 69.8 ($\downarrow 6.6$) 45.7 ($\downarrow 10.3$) 91.0 ($\downarrow 2.5$) 82.9 ($\downarrow 3.6$) 68.4 ($\downarrow 6.4$)
3.3. Optimization Perspective
@@
$\max_{\ell\in\mathcal{L}}\;\;\mathbb{E}_{t\sim\mathcal{T}}\!\left[\mathrm{Success}(t\mid\ell)\right]\;-\;\lambda\,\Omega(\ell),$
@@
where $\mathrm{Success}(t\mid\ell)$ indicates whether $\ell$ enables the system to produce a program that achieves the correct optimal objective for task $t$, and $\Omega(\ell)$ quantifies library complexity (e.g., number of insights or redundancy-adjusted size). Under our problem design—bounded and continuous property of $\text{Success}(\cdot)$ and $\Omega(\cdot)$, sufficient exploration under solver verification, and bounded merging—the refinement dynamics converge to a locally optimal library. In Appendix A, we provide a conceptual sketch that shows that convergence holds: Since refinement in the second phase strictly improves the first term, while verified merging in the first phase reduces the second term without diminishing the first, sufficient exploration combined with iterative cycles of library learning and evolution ensures convergence to a local optimum. Given the inherent ambiguity of natural language and stochasticity in LLM outputs, we present this perspective not as a strict theorem but as a principled justification for the acquisition–refinement design and redundancy-reduction operations.
4. EXPERIMENTS
@@
Our experiments are designed to reflect the requirements that arise in real-world optimization and operations research applications. In these settings, methods are expected not only to perform well on standard benchmarks but also to transfer across domains, remain effective when limited supervision is available, improve steadily as more data become available, and offer results that can be inspected and audited. We organize our empirical evaluation around four questions: (1) How well does the method generalize across domains? (2) How effective are the insight content structure and library refinement mechanisms? (3) Does performance improve consistently as more training data become available? (4) Does the library size converge as training progresses?
4.1. Experimental Setup
@@
Our experiments are conducted on a dataset of 452 problem instances, aggregated from four real-world optimization task datasets—the NLP4LP (AhmadiTeshnizi et al., 2024), NL4OPT (Ramamonjison et al., 2022), IndustryOR (Huang et al., 2025), MAMO (ComplexLP) (Huang et al., 2024). These collections span various formulation types and originate from diverse sources, which include academic papers, textbooks, and real‐world industry scenarios. Detailed descriptions of these datasets are provided in Appendix B.1.
@@
We perform stratified sampling within each dataset, randomly partitioning $70\%$ for training and $30\%$ for testing. The experience library is constructed only from training tasks, which ensures that training-derived insight examples do not leak into the test set. To assess out-of-distribution generalization, we also evaluate on LogiOR (Yang et al., 2025a) and Optibench (Yang et al., 2024).
@@
Unless otherwise specified, gpt-4o (OpenAI, 2024) with temperature 0 is used as the backbone. We use the success rate as the primary evaluation metric, following the evaluation protocol of Yang et al. (2025a), whereby a task is considered successful if the LLM-generated optimal value closely aligns with the provided ground-truth solution. We also report micro- and macro-averaged accuracy for both test-split and out-of-distribution datasets (see Appendix B.2 for definitions).
Baselines. We evaluate against three families of baselines. (i) Prompt Designing: directly generate the mathematical model from manually crafted prompts, which include OptiMUS (AhmadiTeshnizi et al., 2024), and ORThought (Yang et al., 2025a). (ii) Fine-tuning: adapt model parameters through fine-tuning, including ORLM (Huang et al., 2025), built on LLaMa3-8B, and LLMOPT (Jiang et al., 2025), built on Qwen2.5-14B (the latest open-source version available after their paper). (iii) Experience Learning: Accumulate reusable knowledge from past task-solving attempts to improve decision quality across tasks, including two representative works Reflexion (Shinn et al., 2023) and ExpeL (Zhao et al., 2024). For ExpeL, we adapt the original framework to operations research tasks by making minor changes to task-related prompts and evaluation criteria.
4.2. Overall Performance